# Machine learning classification of black holes in the mass–spin diagram

### v1

@@ V Discussion
These results have multiple implications. First, the $m-\chi$ diagram provides a unifying visual and statistical framework for classifying BHs across all mass scales, allowing direct comparison of theory and data in a physically meaningful space. Second, the success of VAEs trained on raw black hole data highlights the power of representation learning in gravitational-wave astrophysics, especially for population classification and source characterization. Third, the ability to distinguish spin prescriptions and formation channels for population-level inference is an active field of research with rich machine learning applicability, where current simple usages of the mass–spin diagram can be extended.
Looking forward, this framework can be improved by incorporating merger products, eccentricity, environmental gas effects, and host galaxy priors into the mass–spin space. Moreover, combining our unsupervised embeddings with supervised classification trained on population synthesis outputs may yield hybrid models capable of probabilistically assigning formation channels to individual detections. As gravitational-wave catalogs grow in size and fidelity, these methods will become increasingly important to the challenging task of decoding the astrophysical formation histories fossilized in the mass and spin distributions of black holes.
In the HR diagram, stars on the main sequence occupy the primary band defined by core hydrogen burning in hydrostatic equilibrium, a fundamental and long-lived phase. We suggest that analogue for black holes may be tied to mass–spin growth pathways, such as prolonged, coherent accretion episodes that define a trajectory “Cosmic Accretion” as the black hole evolves in redshift.
Numerous theoretical and cosmological simulation studies support the universal nature of accretion-driven mass and spin growth [121, 122, 123, 124]. Generally, such simulations find that gas accretion dominates supermassive black hole growth in the high-redshift universe, allowing for efficient spin-up, while galaxy mergers dominate at lower redshifts, leading to more random spin orientations and reduced spin magnitudes; together revealing spin–mass trends for intermediate and supermassive black holes. A recent study found a three-phase evolution [125]: initial rapid accretion spin-up, a chaotic early epoch for moderate-mass black holes, and a final phase dominated by coherent accretion and spin alignment via the Bardeen-Petterson effect, implying the existence of a growth pathway that creates a persistent mass–spin ladder in the mass–spin diagram. A recent general relativistic magneto-hydrodynamic simulation showed the interplay of accretion physics and observable spin precession effects in gravitational-wave signals [126].
@@
The Cosmic Accretion main sequence serves as a locus in mass–spin space determined by well-understood general relativistic accretion physics across diverse formation environments and pathways. Deviations from a track, such as low-spin natal black holes, hierarchical merger remnants, or chaotic accretors, play the role of post-main-sequence or blue-loop outliers. Black holes (and binaries) in isolation will be static on the diagram unless for when they merger, resulting in a merger-tree structure analogue to the main sequence (i.e., can there be multiple black hole main sequences?). This hierarchy helps structure the entire mass–spin plane as black holes steadily move along the Cosmic Accretion track. Meanwhile, GW-detected, merger-product BHs could constitute another primary cluster(s) interpretable as another parallel main sequence to Cosmic Accretion, i.e. black holes appear and disappear as merger products or remnants, and the two are ultimately connected via cosmological expansion history.
@@
The future of GW astronomy is bright with the construction and planning of many detectors across the GW frequency spectrum, with SKA [127] and DAS-2000 [128] in the nano-Hz regime, LISA [64], astrometric facilities such as Gaia [129] and Theia [130] in the micro-Hz regime, TianQin [131], and Taiji [132] in the milli-Hz regime, and Cosmic Explorer [119] and Einstein Telescope [120] in the Hz to kilo-Hz regime. Interestingly, one can perform similar analysis as done here on a diagram defined by GW-frequency and black hole spin to compare the spins of binary mergers as a function of GW frequency. Such multi-banding studies could compare the populations detectable by pulsar timing arrays and LISA (supermassive black holes) with those detectable by LVK, Cosmic Explorer and Einstein Telescope (stellar-mass black holes), by their spin evolution and with both sets of detectors peering into the intermediate-mass regime. Ultimately, the potential for multi-messenger applications of the mass–spin diagram are also plentiful with the development of future X-ray facilities such as AXIS [133] which will provide unprecedented measurements of accreting systems. The classification of joint EM and GW populations, as outlined e.g. in Fig. 6, will be an important direction of future work.
@@ Acknowledgements.
The Authors thank Nicole Ndegwa for support, inspiration, and perspicacious suggestions for visualizations. We also thank Michael Kesden for inspiring early comments and Austin MacMaster for insightful discussions. N.S. and S.S.H. are supported by the Natural Sciences and Engineering Research Council of Canada through the Canada Research Chairs and Discovery Grants programs. This research was enabled by use of the GREX cluster (https://um-grex.github.io/grex-docs/friends/alliancecan/) and the Digital Research Alliance of Canada (alliancecan.ca). Simulations in this paper made use of PyTorch [39], Scikit-learn [40], SciPy [112], NumPy [134], Pandas [135, 136], h5py [137], matplotlib [138], and the COMPAS rapid binary population synthesis code (version 02.42.01) which is freely available at http://github.com/TeamCOMPAS/COMPAS.

### v2

@@ V Discussion
Out results have multiple implications. First the mass–spin diagram provides a visual and statistical framework for classifying BHs across all mass scales, allowing direct comparison of theory and data in a physically meaningful space. Machine learning of gravitational-wave populations in the diagram is useful for studying formation possibilities. This framework can be improved by incorporating merger products, eccentricity, environmental gas effects, and host galaxy priors into the mass–spin space. As observed catalogs grow in size and fidelity, these methods will become increasingly useful for inferring formation histories of black holes. In the HR diagram, stars on the main sequence occupy the primary band defined by core hydrogen burning in hydrostatic equilibrium, a fundamental and long-lived phase. We suggest that analogues for black holes may be tied to mass–spin growth tracks, such as prolonged, coherent accretion episodes, as the black hole evolves in redshift. Numerous theoretical and cosmological simulation studies support the universal nature of accretion-driven mass and spin growth [123, 124, 125, 126]. These simulations imply that gas accretion dominates supermassive black hole growth in the high-redshift universe allowing for efficient spin-up, while galaxy mergers dominate at lower redshifts, possibly providing random spin orientations relative to the host galaxy; together, these create mass–spin tracks for intermediate and supermassive black holes. A recent study found a three-phase evolution, depending on the mass and redshift of the black hole, with predictions for their spin magnitudes and orientations relative to their host galaxies [127]. General relativistic magneto-hydrodynamic simulations show the interplay of accretion physics and generic spin evolution in gravitational-wave events [128].
@@
A “Cosmic Accretion” main sequence serves as a locus in mass–spin space determined by accretion physics across diverse formation environments and pathways. Deviations from a track, such as low-spin natal black holes, hierarchical merger remnants, or chaotic accretors, can play the role of post-main-sequence or blue-loop outliers. Black hole binaries in isolation will be static on the diagram unless they merge, resulting in a merger-tree structure analogue to the main sequence, where black holes appear and disappear as merger products or remnants. Ultimately, the many main sequences would be connected through cosmological evolution of the Universe.
@@
The future of GW astronomy is bright with the construction and planning of many detectors across the GW frequency spectrum, with the Square Kilometer array [129], Meerkat [130], and DAS-2000 [131] in the nano-Hz regime, astrometric facilities such as Theia [132] in the micro-Hz regime, LISA [65], TianQin [133], and Taiji [134] in the milli-Hz regime, and Cosmic Explorers [121] and the Einstein Telescope [122] in the Hz to kilo-Hz regime. Interestingly, one can perform similar analysis as done here on a diagram defined by GW-frequency and black hole spin to compare the spins of binary mergers as a function of GW frequency. Multi-band studies can compare the populations detectable by pulsar timing arrays and LISA (supermassive black holes) with those detectable by LVK, Cosmic Explorer and Einstein Telescope (stellar-mass black holes), by their spin evolution and with both sets of detectors peering into the intermediate-mass black hole regime. Ultimately, the potential for multi-messenger applications of the mass–spin diagram are also plentiful with the development of future X-ray facilities such as AXIS [135] which will provide unprecedented measurements of accreting systems. The classification of joint EM and GW populations, as done in Fig. 6, will be an important direction of future work.
@@ Acknowledgements.
The Authors thank Nicole Ndegwa for support, inspiration, and perspicacious suggestions for visualizations. We also thank Davide Gerosa for helpful feedback on machine learning methods, Michael Kesden for inspiring early comments, and Austin MacMaster for insightful discussions. N.S. and S.S.H. are supported by the Natural Sciences and Engineering Research Council of Canada through the Canada Research Chairs and Discovery Grants programs. This research was enabled by use of the GREX cluster (https://um-grex.github.io/grex-docs/friends/alliancecan/) and the Digital Research Alliance of Canada (alliancecan.ca). Simulations in this paper made use of PyTorch [40], Scikit-learn [41], SciPy [113], NumPy [136], Pandas [137, 138], h5py [139], matplotlib [140], and the COMPAS rapid binary population synthesis code (version 02.42.01) which is freely available at http://github.com/TeamCOMPAS/COMPAS.